OpenStudy (anonymous):

Solve the triangle. A = 32°, a = 19, b = 12

OpenStudy (campbell_st):

so you can find the size of angle B by using the law of sines \[\frac{\sin(32)}{19}=\frac{\sin(B)}{12}\] just solve for B when you have B you can find C by angle sum of a triangle to find the length of side c, use the law of sines again hope this helps

OpenStudy (anonymous):

Why do I do (sin(32)/19) rather than (19/sin(32))?

OpenStudy (campbell_st):

you can... the law of sines works either way but the way I've written it, your unknown, B, is in the numerator

OpenStudy (campbell_st):

you will need to do it that way to find the side c